# Machine Learning, ICML

### v1

@@ 4.1 Experimental setup
For pre-training the neural network, we use CIFAR-10, CIFAR-100 (Krizhevsky & Hinton, 2009), TinyImageNet (Le & Yang, 2015), ImageNet (Deng et al., 2009), Office-31 (Saenko et al., 2010), Office-Home (Venkateswara et al., 2017), Camelyon17-WILDS (Koh et al., 2021), and BREEDS (Santurkar et al., 2020) which leverages class hierarchy of ImageNet (Deng et al., 2009) to create 4 datasets including Living-17, Nonliving-26, Entity-13 and Entity-30. In particular, to avoid time-consuming training, we directly utilize publicly available models pre-trained on Imagenet. Office-31 (Saenko et al., 2010) is collected from 3 different domains: Amazon, DSLR, and Webcam, while Office-Home (Venkateswara et al., 2017) consists of 4 domains: Artistic, Clip Art, Product and Real-World. For these two datasets, we train a neural network on every domain, which means we have 3 pre-trained models for Office-31 and 4 pre-trained models for Office-Home.
@@
In our comprehensive evaluation, we consider 11 datasets with 3 types of distribution shifts: synthetic, natural, and novel subpopulation shift. To verify the effectiveness of our method under the synthetic shift, we use CIFAR-10C, CIFAR-100C, and ImageNet-C (Hendrycks & Dietterich, 2019) that span 19 types of corruption across 5 severity levels, as well as TinyImageNet-C (Hendrycks & Dietterich, 2019) with 15 types of corruption and 5 severity levels. For the natural shift, we use the domains excluded from training from Office-31, Office-Home, and Camelyon17-WILDS as the OOD datasets. For the novel subpopulation shift, we consider the BREEDS benchmark which Living-17, Nonliving-26, Entity-13 and Entity-30 are constructed from ImageNet-C.
@@
To show the versatility of our approach across different architectures, we perform all our experiments on ResNet18, ResNet50 (He et al., 2016) and WRN-50-2 (Zagoruyko & Komodakis, 2016) models. We train them for 20 epochs for CIFAR-10 (Krizhevsky & Hinton, 2009) and for 50 epochs for the other datasets. In all cases, we use SGD with a learning rate of $10^{-3}$, cosine learning rate decay (Loshchilov & Hutter, 2016), a momentum of 0.9, and a batch size of 128.
@@
We measure the performance of all competing methods using the coefficients of determination ($R^{2}$) and Spearman correlation coefficients ($\rho$) calculated between the baseline scores and the true OOD error. To compare the computational efficiency with two self-training methods, we calculate the average evaluation time needed for every test dataset.
Baseline methods.
We consider 9 baselines commonly evaluated in the unsupervised accuracy estimation studies, notably: Rotation Prediction (Rotation) (Deng et al., 2021), Averaged Confidence (ConfScore) (Hendrycks & Gimpel, 2016), Entropy (Guillory et al., 2021), Agreement Score (AgreeScore) (Jiang et al., 2021), Averaged Threshold Confidence (ATC) (Garg et al., 2022), AutoEval (Fréchet) (Deng & Zheng, 2021), Dispersion Score (Dispersion) (Xie et al., 2023), Nuclear Norm (Deng et al., 2023) and ProjNorm (Yu et al., 2022b), where the first seven methods are training-free and the last method is an instance of self-training approaches. In general, Rotation and Agreescore address the problem by constructing an unsupervised loss. ConfScore, Entropy, ATC and Nuclear Norm estimate the test accuracy using the model predictions. Fréchet and ProjNorm gauge the distribution discrepancy as the test accuracy estimation from the feature-representation level and the parameter level, and Dispersion measures the separability of the test feature representation. More details are introduced in Appendix H, and the results compared with Nuclear Norm are discussed in Appendix C.
4.2 Main take-aways
@@
In Table 1, we present the OOD error estimation performance on 11 benchmark datasets across 3 model architectures as measured by $R^{2}$ and $\rho$. We observe that GdScore outperforms existing methods under diverse distribution shifts. Our method achieves an average $R^{2}$ higher than 0.99 on CIFAR-100, while the average $R^{2}$ of the other baseline methods is all below that value. In addition, our method performs stably across different distribution shifts compared with the other existing algorithms. For example, Rotation performs well under the natural shift but experiences a dramatic performance drop under the synthetic shift, ranking from the second best to the eighth. However, our method achieves consistently high performance, ranking the best on average across the three types of distribution shifts. Furthermore, we also provide the visualization of estimation performance in Fig. 1, where we present the scatter plots for Dispersion Score, ProjNorm and GdScore on Entity-13 with ResNet18. From this figure, we find that GdScore and test accuracy perform a strong linear relationship, while the other state-of-the-art methods fail especially when the test error is high. This phenomenon clearly demonstrates the superiority of GdScore in unsupervised accuracy estimation. In the next paragraph, we also demonstrate the computational efficiency of our approach.
@@
In the list of baselines, both our method and ProjNorm (Yu et al., 2022b) belong to self-training methods (Amini et al., 2023). The latter, however, requires costly iterative training on the neural network during evaluation to obtain the complete set of fine-tuned parameters to calculate the distribution discrepancy in network parameters. Compared with ProjNorm, our method only trains the model for one epoch and collects the gradients of the linear classification layer to calculate the gradient norm, which is much more computationally efficient. Fig. 2 presents the comparison of computational efficiency between the two methods on 7 datasets with ResNet50. From this figure, we can see that our method is up to 80% faster than ProjNorm on average. This difference is striking on Office-31 dataset, where our method is not only two orders of magnitude faster than ProjNorm but also improves the $R^{2}$ score by a factor of 2.
@@ 5 Ablation study
Our method involves several algorithmic choices such as the proposed pseudo-labeling strategy, the loss used for backpropagation as well as the number of gradient steps, and the $p$ values of the norm used. Below, we provide an ablation study related to all these choices.
@@
In Fig. LABEL:fig:label_generation, we conduct an ablation study to verify the effectiveness of our label generation strategy by comparing it with ground-truth labels, average labels, full random labels and full pseudo-labeling. From the figure, we observe that while comparable with the ground truth under the synthetic drift and the subpopulation shift, the other strategies lead to a drastic drop in performance under the natural shift. This phenomenon is possibly caused by imprecise gradient calculation based on incorrect pseudo labels. Our labeling strategy performs better on average suggesting that random labels for low-confidence samples provide certain robustness of the score under natural shift.
@@
To demonstrate the impact of different losses on the OOD estimation performance, we compare the standard cross-entropy loss used in our method with the entropy loss for samples with low confidence. Its detailed introduction is in Appendix D. On the other hand, since label smoothing is a simple yet efficient tool for classification performance improvement and model calibration (Müller et al., 2019), we also conduct an ablation study to verify the effectiveness of this regularization by setting the smoothing rate as 0.4. Fig. LABEL:fig:loss_selection illustrates this comparison revealing that standard cross-entropy loss is the most robust choice across different types of distribution shifts. We also note that the entropy loss enhances the estimation performance under the synthetic and the novel subpopulation shifts, but struggles under the natural shift. On the contrary, the label smoothing regularization can increase the performance under the natural shift but decreases it under the synthetic and the novel subpopulation shift.
@@
To illustrate the effect of the norm choice (i.e., $p$ in $L_{p}$-norm) on OOD error estimation, we conduct a sensitivity analysis on three datasets with ResNet18 and gain Fig. LABEL:fig:norm_types. From this figure, we find that there is an obvious decreasing trend as $p$ becomes larger. Especially, when $p$ is smaller than 1, the estimation performance is able to fluctuate within a satisfying range. This is probably because a smaller $p$ (i.e., $10<p<$ 1) makes the $L_{p}$-norm be more suitable for the high-dimensional space, while the $L_{p}$-norm with $p\geq 1$ is likely to ignore gradients that are close to 0 (Wang et al., 2016; Huang et al., 2021). So in this paper, we set $p$ as 0.3 for all datasets and model structures. The following remark provides some theoretical insights into the $L_{p}$-norm of the gradient for $10<p<$ 1. The proof is deferred in Appendix I.4
@@
Let $\mathbf{c}$ be the classifier obtained from $\bm{\omega}_{s}$ after one gradient descent step, i.e., $\mathbf{c}=\bm{\omega}_{s}-\eta\cdot\nabla\mathcal{L}_{T}(\bm{\omega}_{s})$ with $\eta\geq 0$. For any $p\in(0,1)$, we have that
@@
In this part, we aim to understand whether backpropagating through other layers in the neural network can provide a better OOD error estimation. For this, we separate the feature extractor $f_{\mathbf{g}}$ into 3 blocks of layers with roughly equal size and calculate the GrdNorm scores for each of them. Additionally, we also try to gather the gradients over the whole network. Fig. LABEL:fig:layer_selection plots the obtained results on 4 datasets, suggesting that the last layer provides sufficient information to predict the true OOD error.
@@
Here, we train the neural network for $r$ epochs, where $r\in\{1,2,5,10,15,20\}$, and store the gradient vectors of the classification layer for each value of $r$. Fig. LABEL:fig:epochs suggests that the gradient norms after 1 step are sufficient to predict the model performance under distribution shifts. Further training gradually degrades the performance with the increasing $r$. The reason behind the phenomenon is that the gradients in the first epoch contain the most abundant information about the training dataset, while the neural network fine-tuned on the test dataset for several epochs is going to forget previous training categories (Kemker et al., 2018).
@@ 6 Conclusion
In this paper, we demonstrated a strong linear relationship between the magnitude of gradients and the model performance under distribution shifts. We proposed GrdNorm Score, a simple yet efficient method to estimate the ground-truth OOD error by measuring the gradient norm of the last classification layer. Our method can achieve consistently better performance across different distribution shifts than previous works with high computational efficiency. Furthermore, it does not require the detailed architecture of the feature extractors and training datasets. Those properties guarantee that our method can be easily deployed in the real world, and meet practical demands, such as large models and confidential information. We hope that our research throws light on the future exploration of gradient norm for OOD error estimation.
Appendix A Pseudo code of GdScore
@@
Appendix B Choice of proper threshold $\tau$
Threshold 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
CIFAR-10C 0.963 0.963 0.964 0.965 0.971 0.972 0.967 0.962 0.963 0.959
Office-31 0.495 0.498 0.532 0.674 0.685 0.667 0.545 0.451 0.114 0.131
Appendix C Results compared with Nuclear Norm
Here, we compare our method to Nuclear Norm (Deng et al., 2023) across 3 types of distribution shifts with ResNet18, ResNet50 and WRN-50-. Results are shown in Table 3. From this table, we observe our method outperforms Nuclear Norm under synthetic shift and natural shift.
$R^{2}$ $\rho$ $R^{2}$ $\rho$ $R^{2}$ $\rho$ $R^{2}$ $\rho$ $R^{2}$ $\rho$ $R^{2}$ $\rho$
Appendix D Entropy loss for low-confidence samples
Appendix E Related Work
Unsupervised accuracy estimation. Unsupervised accuracy estimation is a vital topic in practical applications due to frequent distribution shifts and infeasible ground-truth labels of the test samples. To comprehensively understand this field, we introduce two main existing settings which are related to this topic.
  1. 1.
    Some works aim to estimate the test accuracy or gauge the accuracy discrepancy between the training and the test set only via the training data (Corneanu et al., 2020; Jiang et al., 2019; Neyshabur et al., 2017; Unterthiner et al., 2020; Yak et al., 2019; Martin & Mahoney, 2020). For example, the model-architecture-based algorithm (Corneanu et al., 2020) derives plenty of persistent topology properties from the training data, which can identify when the model learns to generalize to unseen datasets. However, those algorithms are deployed under the assumption that the training and the test data are drawn from the same distribution, which means they are vulnerable to distribution shifts.
  2. 2.
    Our work belongs to the second setting, which aims to estimate the classification accuracy of a specific test dataset during evaluation using unlabeled test samples and/or labeled training datasets. The main research direction is to explore the negative relationship between the distribution discrepancy and model performance from the space of features (Deng & Zheng, 2021), parameters (Yu et al., 2022b) and labels (Lu et al., 2023). Another popular direction is to design an estimation score via the softmax outputs of the test samples (Guillory et al., 2021; Jiang et al., 2021; Guillory et al., 2021; Garg et al., 2022), which heavily relies on model calibration. Some works also learn from the field of unsupervised learning, such as agreement across multiple classifiers (Jiang et al., 2021; Madani et al., 2004; Platanios et al., 2016, 2017) and image rotation (Deng et al., 2021). In addition, the property of the test datasets presented during evaluation has been also studied recently (Xie et al., 2023). To the best of our knowledge, our work is the first to study the linear relationship between the gradients and model performance.
Gradients in generalization. The role of gradients in generalization has attracted increasing attention recently. To gauges the generalization performance of the hypothesis learned from the training data on unseen samples, known as the out-of-sample error (Hardt et al., 2016; London, 2017; Rivasplata et al., 2018), many studies try to provide a tight upper bound for generalization error from the view of gradient descent theoretically, indicating that gradients correlate with the discrepancy between the empirical loss and the population loss (Li et al., 2019; Chatterjee, 2020; Negrea et al., 2019; An et al., 2020). However, those works assume that seen to unseen data are from the identical distribution, while unsupervised accuracy estimation discusses a more complex and realistic issue that they come from different distributions. Under distribution shift, gradients are also explored. For example, Mansilla et al. (2021) clips the conflicting gradients emerging in domain shift scenarios and promote gradient agreement across multiple domains via the gradient agreement strategy. Similarly, Zhao et al. (2022) designs a gradient-based regularization term to make the optimizer find the flat minima. In out-of-distribution (OOD) detection which goal is to determine whether a given sample is in-distribution (ID) or out-of-distribution (Hendrycks & Gimpel, 2016; Hendrycks et al., 2018; Liu et al., 2020; Yang et al., 2021; Liang et al., 2017), (Huang et al., 2021) finds that ID data usually have higher gradient magnitude than OOD data from current source distribution to an uniform distribution. However, despite their empirical success, the relationship between gradients and generalization is still unclear.
Appendix F Discussion: relation to Huang et al. (2021)
A current work, GradNorm (Huang et al., 2021), employs gradients to detect OOD samples which labels belong to a different label space from the training data. It gauges the magnitude of gradients in the classification layer, backpropagated from a KL-divergence between the softmax probability and a uniform distribution. Compared with GradNorm, our method bears three critical differences, in terms of the problem setting, methodology, and theoretical insights. We also empirically demonstrate the superiority of our method in Table 5. 1) Problem setting: GradNorm focuses on OOD detection, which aims to determine whether a given sample is in-distribution (ID) or out-of-distribution (Hendrycks & Gimpel, 2016; Hendrycks et al., 2018; Liu et al., 2020; Yang et al., 2021; Liang et al., 2017), while our method aims to estimate the test accuracy without ground-truth test labels. It requires GradNorm should be an instance-level score for classification, but our method is a dataset-level score for linear regression. Furthermore, in OOD detection, the label spaces of OOD data and training data are disjoint, while in unsupervised test accuracy, the training and the OOD label spaces are shared. The two are also evaluated differently: AUROC score for OOD detection and correlation coefficients for error estimation. Those differences are summarized in Table 4.
Learning Problem Goal Scope Metric
OOD detection Predict ID/OOD $\tilde{\mathbf{x}_{i}}$ AUROC
2) Methodology: GradNorm obtains the magnitude of gradients via a KL-divergence loss measuring the distribution distance from the training distribution to an uniform distribution, while our method obtain them using a standard cross entropy loss with the specifically-designed label generation strategy. GradNorm assumes that OOD data should have lower magnitude of gradients, but in our cases, test data are certified to have higher magnitude of gradients. Table 5 presents the performance comparison of the two methods in unsupervised accuracy estimation on 7 datasets across 3 types of distribution shifts with ResNet18. It illustrates that GradNorm is inferior to our approach for unsupervised accuracy estimation suggesting that the two problems cannot be tackled with the same tools. 3) Theoretical insights: GradNorm is certified to captures the joint information between features and outputs to detect OOD data from the oncoming dataset. However, the reason why OOD data have lower magnitude of gradients is still unclear. Our method provides a more clear and direct theoretical insights to demonstrate the relationship between gradients and test accuracy even under distribution shift, which further inspires future work to address generalization issues from the view of gradients
(Huang et al., 2021) 0.951 0.978 0.894 0.596 0.848 0.964 0.942
Ours 0.972 0.983 0.971 0.675 0.876 0.970 0.949
Appendix G Influence of Calibration Error
We have mentioned earlier that, in theory, the proposed pseudo-labeling strategy depends on how well the prediction probabilities are calibrated. In degraded cases, this can have a negative impact on our approach, e.g., one can imagine a model that outputs only one-hot probabilities with not a high accuracy. However, this it is generally not the case. Indeed, in practice, we do not need to have a perfectly calibrated model as we employ a mixed strategy that assigns pseudo-labels to high-confidence examples and random labels to low-confidence ones. The recent success of applying self-training models to different problems (Sohn et al., 2020; Dong et al., 2021; Yu et al., 2022a) provides evidence of the suitability of the label generation strategy we adopted. When we speak of deep neural networks, which are widely accepted to be poorly calibrated, Minderer et al. (2021) showed that modern SOTA image models tend to be well-calibrated across distribution shifts. To demonstrate it empirically, in Table 6, we provide the expected calibration error (ECE, (Guo et al., 2017)) of ResNet18, one of the considered base models, depending on a difficulty of test data. For this, we test first on CIFAR-10 (ID), and then on CIFAR-10C corrupted by brightness across diverse severity from 1 to 5. We can see that ECE is very low for ID data and remains relatively low across all levels of corruption severity, which shows that ResNet is quite well-calibrated on CIFAR-10.
Corruption Severity ID 1 2 3 4 5
ECE 0.0067 0.0223 0.0230 0.0243 0.0255 0.0339
On the other hand, in the case of more complex distribution shift like Office-31 data set, we can see that the calibration error has been increased noticeably (Table 7). It is interesting to analyze this result together with Figure LABEL:fig:label_generation of the main paper, where we compared the results between the usual pseudo-labeling strategy and the proposed one. Although our method has room for improvement compared to the oracle method, it is also significantly better than ”pseudo-labels”, indicating that the proposed label generation strategy is less sensitive to the calibration error.
Domain DSLR (ID) Amazon Webcam
ECE 0.2183 0.2167 0.4408
Appendix H Baseline methods

### v3

@@ 5.1 Experimental setup
For pre-training the neural network, we use CIFAR-10, CIFAR-100 (Krizhevsky & Hinton, 2009), TinyImageNet (Le & Yang, 2015), ImageNet (Deng et al., 2009), Office-31 (Saenko et al., 2010), Office-Home (Venkateswara et al., 2017), Camelyon17-WILDS (Koh et al., 2021), and BREEDS (Santurkar et al., 2020) which leverages class hierarchy of ImageNet (Deng et al., 2009) to create 4 datasets including Living-17, Nonliving-26, Entity-13 and Entity-30. In particular, to avoid time-consuming training, we directly utilize publicly available models pre-trained on Imagenet. For Office-31 and Office-Home, we train a neural network on every domain.
Table 1: Performance comparison on 11 benchmark datasets with ResNet18, ResNet50, and WRN-50-2, where $R^{2}$ refers to coefficients of determination, and $\rho$ refers to the absolute value of Spearman correlation coefficients (higher is better). The best results are highlighted in bold.
@@
In our comprehensive evaluation, we consider 11 datasets with 3 types of distribution shifts: synthetic, natural, and novel subpopulation shift. To verify the effectiveness of our method under the synthetic shift, we use CIFAR-10C, CIFAR-100C, and ImageNet-C (Hendrycks & Dietterich, 2019) that span 19 types of corruption across 5 severity levels, as well as TinyImageNet-C (Hendrycks & Dietterich, 2019) with 15 types of corruption and 5 severity levels. For the natural shift, we use the domains excluded from training from Office-31, Office-Home, and Camelyon17-WILDS as the OOD datasets. For the novel subpopulation shift, we consider the BREEDS benchmarks, namely, Living-17, Nonliving-26, Entity-13, and Entity-30, which are constructed from ImageNet-C.
@@
To show the versatility of our approach across different architectures, we perform all our experiments on ResNet18, ResNet50 (He et al., 2016) and WRN-50-2 (Zagoruyko & Komodakis, 2016) models. We train them for 20 epochs for CIFAR-10 (Krizhevsky & Hinton, 2009) and 50 epochs for the other datasets. In all cases, we use SGD with a learning rate of $10^{-3}$, cosine learning rate decay (Loshchilov & Hutter, 2016), a momentum of 0.9, and a batch size of 128. For all experiments, we use $p=0.3$ to compute GdScore.
@@
We measure the performance of all competing methods using the coefficients of determination ($R^{2}$) and Spearman correlation coefficients ($\rho$) calculated between the baseline scores and the true test error. To compare the computational efficiency with two self-training methods, we calculate the average evaluation time needed for every test dataset.
Baselines.
We compare our method GdScore with $8$ baselines commonly considered in the unsupervised accuracy estimation literature: Rotation Prediction (Rotation) (Deng et al., 2021), Averaged Confidence (ConfScore) (Hendrycks & Gimpel, 2016), Entropy (Guillory et al., 2021), Agreement Score (AgreeScore) (Jiang et al., 2021), Averaged Threshold Confidence (ATC) (Garg et al., 2022), AutoEval (Fréchet) (Deng & Zheng, 2021), Dispersion Score (Dispersion) (Xie et al., 2023), and ProjNorm (Yu et al., 2022b). The first six methods are training-free and the last method is an instance of self-training approaches. More details about the baselines can be found in Appendix C.
5.2 Main takeaways
@@
In Table 1, we present the OOD error estimation performance on 11 benchmark datasets across 3 model architectures as measured by $R^{2}$ and $\rho$. We observe that GdScore outperforms existing methods under diverse distribution shifts. Our method achieves an average $R^{2}$ higher than $0.99$ on CIFAR-100, while the average $R^{2}$ of the other baselines is always below. In addition, our method performs stably across different distribution shifts compared with the other existing algorithms. For example, Rotation performs well under the natural shift but experiences a dramatic performance drop under the synthetic shift, ranking from the second best to the eighth. However, our method achieves consistently high performance, ranking the best on average across the three types of distribution shifts.
(a) Label generation strategies
(b) Loss selection
Figure 3: Performance comparison ($R^{2}$) on $7$ datasets with ResNet18 between (a) different label generation strategies and (b) different types of losses across $3$ types of distribution shifts. Results confirm that our proposed method performs better on average across various datasets and types of shifts.
Furthermore, we provide the visualization of estimation performance in Fig. 1, where we present the scatter plots for Dispersion Score, ProjNorm and GdScore on Entity-13 with ResNet18. We can see that GdScore and test accuracy have a strong linear relationship, while the other state-of-the-art methods struggle to have a linear correlation in cases when the test error is high. This phenomenon demonstrates the superiority of GdScore in unsupervised accuracy estimation. In the next paragraph, we also demonstrate the computational efficiency of our approach.
@@
In the list of baselines, both our method and ProjNorm (Yu et al., 2022b) belong to self-training methods (Amini et al., 2023). The latter, however, requires costly iterative training on the neural network during evaluation to obtain the complete set of fine-tuned parameters to calculate the distribution discrepancy in network parameters. Compared with ProjNorm, our method only trains the model for one epoch and collects the gradients of the linear classification layer to calculate the gradient norm, which is much more computationally efficient. Fig. 2 presents the comparison of computational efficiency between the two methods on 7 datasets with ResNet50. From this figure, we can see that our method is up to 80% faster than ProjNorm on average. This difference is striking on the Office-31 dataset, where our method is not only two orders of magnitude faster than ProjNorm but also improves the $R^{2}$ score by a factor of 2.
Figure 4: Robustness comparison for all estimation baselines across diverse distribution shifts with ResNet18.
Robustness of our approach.
GdScore achieves great performance across all datasets, architectures, and types of shifts (Table 1). To highlight the robustness of our approach, we compare the distributions of $R^{2}$ in Figure 4, including as additional baseline the very recent Nuclear Norm (Deng et al., 2023) (more results in Appendix F). We show that GdScore is the best and most stable approach on $10$ datasets (except ImageNet).
@@ 6 Ablation study
In this section, we will conduct comprehensive ablation studies to verify the efficiency of our proposed label-generation strategy, and choice of hyperparameters such as $\tau$ and $p$ as well as self-training settings to obtain GdScore such as epochs, blocks, and losses.
@@
In Fig. 3(a), we conduct an ablation study to verify the effectiveness of our label generation strategy by comparing it with ground-truth labels, uniform labels (Huang et al., 2021), full random labels, and full pseudo-labeling. From the figure, we observe that while comparable with the ground truth under the synthetic drift and the subpopulation shift, the other strategies lead to a drastic drop in performance under the natural shift. This phenomenon is possibly caused by imprecise gradient calculation based on incorrect pseudo labels. Our labeling strategy performs better on average suggesting that random labels for low-confidence samples provide certain robustness of the score under natural shift.
(a) Norm types
(b) Layer selection
(c) Epoch selection
Figure 5: Sensitivity analysis on the effect of (a) norm types, (b) layer selection for gradients, and (c) epoch selection. The first and the third experiments are conducted on CIFAR-10C, CIFAR-100C, and TinyImageNet-C, while the second experiment includes TinyImageNet-C, Office-31, Office-Home, and WILDS-FMoV (Koh et al., 2021). All experiments are conducted with ResNet18.
@@
To demonstrate the impact of different losses on the test accuracy estimation performance, we compare the standard cross-entropy loss used in our method with the entropy loss for samples with low confidence (see definition in Appendix D). Moreover, we verify the effectiveness of the label smoothing, a simple yet effective tool for model calibration and performance improvement (Müller et al., 2019), by setting the smoothing rate as 0.4.
Fig. 3(b) illustrates this comparison revealing that standard cross-entropy loss is the most robust choice across different types of distribution shifts. We also note that the entropy loss enhances the estimation performance under the synthetic and the novel subpopulation shifts, but struggles under the natural shift. On the contrary, the label smoothing regularization can increase the performance under the natural shift but decrease it under the synthetic and the novel subpopulation shifts.
@@
To illustrate the effect of the choice of $L_{p}$-norm on the performance, we conduct a sensitivity analysis on three datasets with ResNet18 summarized in Figure 5(a). We can see that there is an obvious decreasing trend as $p$ becomes larger. Especially, when $p$ is smaller than 1, the estimation performance can fluctuate within a satisfying range. This is probably because a smaller $p$ (i.e., $10<p<$ 1) makes the $L_{p}$-norm more suitable for the high-dimensional space, while the $L_{p}$-norm with $p\geq 1$ is likely to ignore gradients that are close to 0 (Wang et al., 2016; Huang et al., 2021), so for all experiments, we used $p=0.3$. The remark below, whose proof we defer to Appendix H.4, provides some theoretical insights into the $L_{p}$-norm of the gradient for $10<p<$ 1.
@@
Let $\mathbf{c}$ be the classifier obtained from $\bm{\omega}_{s}$ after one gradient descent step, i.e., $\mathbf{c}=\bm{\omega}_{s}-\eta\cdot\nabla\mathcal{L}_{T}(\bm{\omega}_{s})$ with $\eta\geq 0$. For any $p\in(0,1)$, we have
@@
In this part, we aim to understand whether backpropagation through other layers in the neural network can provide a better test accuracy estimate. For this, we separate the feature extractor $f_{\mathbf{g}}$ into 3 blocks of layers with roughly equal size and calculate the GrdNorm scores for each of them. Additionally, we also try to gather the gradients over the whole network. Fig. 5(b) plots the obtained results on 4 datasets, suggesting that the last layer provides sufficient information to predict the true test accuracy in an unsupervised way.
@@
Here, we train the neural network for $r$ epochs, where $r\in\{1,2,5,10,15,20\}$, and store the gradient vectors of the classification layer for each value of $r$. Fig. 5(c) suggests that the gradient norms after 1 step are sufficient to predict the model performance under distribution shifts. Further training gradually degrades the performance with the increasing $r$. The reason behind the phenomenon is that the gradients in the first epoch contain the most abundant information about the training dataset, while the neural network fine-tuned on the test dataset for several epochs is going to forget previous training categories (Kemker et al., 2018).
Choice of proper threshold $\tau$
Threshold 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
CIFAR-10C 0.963 0.963 0.964 0.965 0.971 0.972 0.967 0.962 0.963 0.959
Office-31 0.495 0.498 0.532 0.674 0.685 0.667 0.545 0.451 0.114 0.131
7 Connection to GradNorm
A current work, GradNorm (Huang et al., 2021), employs gradients to detect OOD samples whose labels belong to a different label space from the training data. It gauges the magnitude of gradients in the classification layer, backpropagated from a KL divergence between the softmax probability and uniform distribution. Compared with GradNorm, our method bears three critical differences, in terms of the problem setting, methodology, and theoretical insights.
1) Problem setting: GradNorm focuses on OOD detection, which aims to determine whether a given sample is in-distribution (ID) or out-of-distribution (Hendrycks & Gimpel, 2016; Hendrycks et al., 2018; Liu et al., 2020; Yang et al., 2021; Liang et al., 2017), while our method aims to estimate the test accuracy without ground-truth test labels. It requires GradNorm should be an instance-level score for classification, but our method is a dataset-level score for linear regression. Furthermore, in OOD detection, the label spaces of OOD data and training data are disjoint, while in unsupervised test accuracy, the training and the OOD label spaces are shared. The two are also evaluated differently: AUROC score for OOD detection and correlation coefficients for error estimation. Those differences are summarized in Table 3.
Learning Problem Goal Scope Metric
OOD detection Predict ID/OOD $\tilde{\mathbf{x}_{i}}$ AUROC
2) Methodology: GradNorm obtains the magnitude of gradients via a KL-divergence loss measuring the distribution distance from the training distribution to a uniform distribution, while our method obtains them using a standard cross entropy loss with the specifically designed label generation strategy. GradNorm assumes that OOD data should have a lower magnitude of gradients, but in our cases, test data are certified to have a higher magnitude of gradients. Table 4 presents the performance comparison of the two methods in unsupervised accuracy estimation on 7 datasets across 3 types of distribution shifts with ResNet18. It illustrates that GradNorm is inferior to our approach for unsupervised accuracy estimation suggesting that the two problems cannot be tackled with the same tools.
3) Theoretical insights: GradNorm is certified to capture the joint information between features and outputs to detect OOD data from the oncoming dataset. However, the reason why OOD data have a lower magnitude of gradients is still unclear. Our method provides clearer theoretical insights to demonstrate the relationship between gradients and test accuracy even under distribution shift, which further inspires future work to address generalization issues from the view of gradients
(Huang et al., 2021) 0.951 0.978 0.894 0.596 0.848 0.964 0.942
Ours 0.972 0.983 0.971 0.675 0.876 0.970 0.949
@@ 8 Conclusion
In this paper, we showcased the strong linear relationship between the magnitude of gradients and the model performance under distribution shifts. We proposed GdScore, a simple yet efficient method to estimate the ground-truth test accuracy error by measuring the gradient magnitude of the last classification layer. Our method consistently achieves superior performance across various distribution shifts than previous works. Furthermore, it does not require the detailed architecture of the feature extractors and training datasets. Those properties guarantee that our method can be easily deployed in the real world, and meet practical demands, such as large models and confidential information. We hope that our research sheds new light on the usefulness of gradient norms for unsupervised accuracy estimation.
Acknowledgments
This research is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISGAward No: AISG2-GC-2023-009).
Appendix A Appendix
Appendix B Pseudo-code of GdScore
@@
Algorithm 1 Unsupervised Accuracy Estimation via GdScore
Appendix C Baselines